X^2+y^2-12x-8y+27=0
Answer by ewatrrr(24785)

(x-6)^2 + (y-4)^2 +27 - 36 - 16 = 0
(x-6)^2 + (y-4)^2 = 25
Circle
C(6,4), r = 5